Please help me with the following question:
Find the slope of a line containing points (-2,8)and(3,5).
Thank you!

Answer by jim_thompson5910(35256)

Use the slope formula

m = (y2-y1)/(x2-x1)


m = (5-8)/(3-(-2))


m = (5-8)/(3+2)


m = -3/5


So the slope of the line that passes through the two points is -3/5